# Integral representation of solutions to initial-boundary value problems in the framework of the hyperbolic heat equation

### v1

@@ A.2 The initial state is non-zero and the boundaries are thermally insulating
Here we assume that the initial state is non-zero and the boundaries are thermally insulating ($\gamma_{0}=\gamma_{l}=0$ and $g=0$, $h=0$). In this case both equations are homogeneous:
$\partial_{t}e+\partial_{x}q=0,\quad x\in(0,l),\quad t>0,$ (A.6a)
$\tau\partial_{t}q+q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0.$ (A.6b)
the initial conditions remain the same:
$\left.e\right|_{t=0}=\varphi(x),\quad\left.q\right|_{t=0}=\psi(x),$ (A.7)
and the boundary conditions take the form
$\left.q\right|_{x=0}=0,\quad\left.q\right|_{x=l}=0.$ (A.8)
The solution to the problem (A.6)–(A.8) is given by
@@
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k,$ (A.9)
where $\hat{\boldsymbol{u}}_{0}(k)$ is given by Eq. (3.7),
@@
$S(k)\boldsymbol{a}^{\prime}(k,t)$ $=\frac{\mathrm{i}k\beta}{\omega_{1}-\omega_{2}}\bigg{[}-\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\tilde{e}^{(1)}(0,t)+\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\tilde{e}^{(2)}(0,t)\bigg{]},$ (A.10a)
@@
$S(k)\boldsymbol{b}^{\prime}(k,t)$ $=\frac{\mathrm{i}k\beta}{\omega_{1}-\omega_{2}}\bigg{[}-\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\tilde{e}^{(1)}(l,t)+\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\tilde{e}^{(2)}(l,t)\bigg{]},$ (A.10b)
(see Eq. (3.29), where $\tilde{g}^{(1,2)}=0$, $\tilde{h}^{(1,2)}=0$ and $\sigma_{0,l}^{(1,2)}(k)=\mathrm{i}k\beta$), the functions $\tilde{e}^{(1,2)}$ are given by
@@
$-\mathrm{i}k\beta\begin{pmatrix}\tilde{e}^{(1)}(0,t)\\[2.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}$ $=\frac{1}{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}\bigg{[}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\\ 1\end{pmatrix}\varPhi_{1}(k)+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\\ 1\end{pmatrix}\varPhi_{1}(-k)\bigg{]},$ (A.11a)
@@
$\mathrm{i}k\beta\begin{pmatrix}\tilde{e}^{(2)}(0,t)\\ \tilde{e}^{(2)}(l,t)\end{pmatrix}$ $=\frac{1}{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}\bigg{[}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\\ 1\end{pmatrix}\varPhi_{2}(k)+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\\ 1\end{pmatrix}\varPhi_{2}(-k)\bigg{]},$ (A.11b)
(see Eq. (3.30), where $V_{1,2}(k,t)=\varPhi_{1,2}(k)$ and $\varDelta_{1,2}(k)=-(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})k^{2}\beta^{2}$).
@@
Deforming the paths $\partial D_{0}^{\pm}$ back to the real axis, we see that the representation (A.9) is transformed into the representation
@@
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k+\frac{\pi\mathrm{i}}{2\pi}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]\\ +\frac{1}{2\pi}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.25pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.29166pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-1.875pt}}\!\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k+\frac{\pi\mathrm{i}}{2\pi}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right],$ (A.12)
where $k_{n}=\pi n/l$ are the poles. It follows from Eqs. (A.10) and (A.11) the relation
$S(k)\boldsymbol{a}^{\prime}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)=\frac{1}{\omega_{1}-\omega_{2}}\begin{pmatrix}\varPhi_{1}(k)+\varPhi_{2}(k)\\[2.0pt] \left[\omega_{1}\varPhi_{1}(k)+\omega_{2}\varPhi_{2}(k)\right]/\mathrm{i}k\end{pmatrix}\\[2.0pt] =\frac{1}{\omega_{1}-\omega_{2}}\left[\begin{pmatrix}(\omega_{1}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t}-\omega_{2}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t})\\[2.0pt] \mathrm{i}k\beta(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t})\end{pmatrix}\hat{\varphi}(k)+\begin{pmatrix}\mathrm{i}k(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t})\\[2.0pt] (\omega_{1}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}t}-\omega_{2}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}t})\end{pmatrix}\hat{\psi}(k)\right]\\[2.0pt] =\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)$
@@
Hence the difference of the two principal-value integrals in Eq. (A.12) is equal to the first integral, and, therefore, the net contribution of the three integrals to the solution is equal to zero. Thus the representation (A.12) takes the form
@@
Using the equalities $\mathop{\mathrm{e}}\nolimits^{\pm\mathrm{i}k_{n}l}=(-1)^{n}$ and $[(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})_{k}^{\prime}]|_{k=k_{n}}=2\mathrm{i}l(-1)^{n}$ we see that the residues are equal to
$\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]=\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right]\\[2.0pt] =\frac{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}k_{n}x}}{2\mathrm{i}l(\omega_{1}^{(n)}-\omega_{2}^{(n)})}\begin{pmatrix}\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})+\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\\[2.0pt] \left\{\omega_{1}^{(n)}\left[\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})\right]+\omega_{2}^{(n)}\left[\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\right]\right\}/\mathrm{i}k_{n}\end{pmatrix},$
where $\omega_{1,2}^{(n)}\equiv\omega_{1,2}(k_{n})$. The functions $\varPhi_{1,2}$ are given by Eq. (3.16), hence
$\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})+\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\\[2.0pt] =(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t})\left[\hat{\varphi}(k_{n})+\hat{\varphi}(-k_{n})\right]\\[2.0pt] +\mathrm{i}k_{n}(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\left[\hat{\psi}(k_{n})-\hat{\psi}(-k_{n})\right]\\[2.0pt] =2\left[(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t})\varphi_{n}+k_{n}(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\psi_{n}\right],\quad n\in\mathbb{Z}\backslash\{0\},$
@@
$\frac{\omega_{1}^{(n)}\left[\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})\right]+\omega_{2}^{(n)}\left[\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\right]}{\mathrm{i}k_{n}}\\[2.0pt] =\mathrm{i}k_{n}\beta(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\left[\hat{\varphi}(k_{n})+\hat{\varphi}(-k_{n})\right]\\[2.0pt] +(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\left[\hat{\psi}(k_{n})-\hat{\psi}(-k_{n})\right]\\[2.0pt] =2\mathrm{i}\left[k_{n}\beta(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\varphi_{n}-(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\psi_{n}\right],\quad n\in\mathbb{Z}\backslash\{0\},$
@@
$\boldsymbol{u}(x,t)\equiv\begin{pmatrix}e(x,t)\\[2.0pt] q(x,t)\end{pmatrix}=\begin{pmatrix}\varphi_{0}\\[3.0pt] 0\end{pmatrix}+\sum_{n=1}^{\infty}\frac{1}{\omega_{1}^{(n)}-\omega_{2}^{(n)}}\times\\[4.0pt] \times\begin{pmatrix}\big{[}(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t})\varphi_{n}+k_{n}(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\psi_{n}\big{]}\cos(k_{n}x)\\[4.0pt] \big{[}-k_{n}\beta(\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\varphi_{n}+(\omega_{1}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}^{(n)}t}-\omega_{2}^{(n)}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}^{(n)}t})\psi_{n}\big{]}\sin(k_{n}x)\end{pmatrix},$
where $k_{n}=\pi n/l$, $\omega_{1,2}^{(n)}\equiv\omega_{1,2}(k_{n})$ (see Eq. (3.1)). This solution can also be easily obtained by the Fourier method (see the solution in Ref. [13]).

### v2

@@ B.2 The initial state is non-zero and the boundaries are thermally insulating
Here we assume that heat sources are absent ($f=0$), the initial state is non-zero, and the boundaries are thermally insulating ($\gamma_{0}=\gamma_{l}=0$ and $g=0$, $h=0$). In this case the equations are homogeneous:
$\partial_{t}e+\partial_{x}q=0,\quad x\in(0,l),\quad t>0,$ (B.6a)
$\tau\partial_{t}q+q-\mu^{2}\partial_{x}^{2}q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0.$ (B.6b)
the initial conditions are
$\left.e\right|_{t=0}=\varphi(x),\quad\left.q\right|_{t=0}=\psi(x),$ (B.7)
and the boundary conditions are
$\left.q\right|_{x=0}=0,\quad\left.q\right|_{x=l}=0.$ (B.8)
The solution to the problem (B.6)–(B.8) is given by
@@
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k.$ (B.9)
where
@@
$S(k)\boldsymbol{a}^{\prime}(k,t)$ $=\frac{1}{\omega_{1}-\omega_{2}}\bigg[-\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\mathrm{i}k(\beta-\theta\omega_{1})\tilde{e}_{0}^{(1)}(0,t)+\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\mathrm{i}k(\beta-\theta\omega_{2})\tilde{e}_{0}^{(2)}(0,t)\bigg],$ (B.10a)
@@
$S(k)\boldsymbol{b}^{\prime}(k,t)$ $=\frac{1}{\omega_{1}-\omega_{2}}\bigg[-\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\mathrm{i}k(\beta-\theta\omega_{1})\tilde{e}_{0}^{(1)}(l,t)+\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\mathrm{i}k(\beta-\theta\omega_{2})\tilde{e}_{0}^{(2)}(l,t)\bigg],$ (B.10b)
and
@@
$-\mathrm{i}k(\beta-\theta\omega_{1})\begin{pmatrix}\tilde{e}_{0}^{(1)}(0,t)\\[3.0pt] \tilde{e}_{0}^{(1)}(l,t)\end{pmatrix}=\frac{1}{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}\Bigg[$ $\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\\[3.0pt] 1\end{pmatrix}V_{1}(k,t)+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\\[3.0pt] 1\end{pmatrix}V_{1}(-k,t)\Bigg],$ (B.11a)
@@
$\mathrm{i}k(\beta-\theta\omega_{2})\begin{pmatrix}\tilde{e}_{0}^{(2)}(0,t)\\[3.0pt] \tilde{e}_{0}^{(2)}(l,t)\end{pmatrix}=\frac{1}{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}\Bigg[$ $\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\\[3.0pt] 1\end{pmatrix}V_{2}(k,t)+\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\\[3.0pt] 1\end{pmatrix}V_{2}(-k,t)\Bigg].$ (B.11b)
@@
Deforming the paths $\partial D_{0}^{\pm}$ back to the real axis, we see that the representation (B.9) is transformed into the representation
@@
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k+\frac{\pi\mathrm{i}}{2\pi}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]\\ +\frac{1}{2\pi}\mathchoice{{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}}{{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}}{{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}}\!\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k+\frac{\pi\mathrm{i}}{2\pi}\sum_{n\in\mathbb{Z}}\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right],$ (B.12)
where $k_{n}=\pi n/l$ are the poles. It follows from Eqs. (B.10) and (B.11) the relation
$S(k)\boldsymbol{a}^{\prime}(k,t)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)=S(k)\boldsymbol{\varPhi}(k)=\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k).$
@@
Hence the difference of the two principal-value integrals in Eq. (B.12) is equal to the first integral, and, therefore, the net contribution of the three integrals to the solution is equal to zero. Thus the representation (B.12) takes the form
@@
Using the equalities $\mathop{\mathrm{e}}\nolimits^{\pm\mathrm{i}k_{n}l}=(-1)^{n}$ and $(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})_{k}^{\prime}]|_{k=k_{n}}=2\mathrm{i}l(-1)^{n}$ we see that the residues are equal to
$\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]=\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right]\\[2.0pt] =\frac{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}k_{n}x}}{2\mathrm{i}l}\left\{S(k_{n})\big[\boldsymbol{\varPhi}(k_{n})+\boldsymbol{\varPhi}(-k_{n})\big]\right\}\\[2.0pt] =\frac{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}k_{n}x}}{2\mathrm{i}l(\omega_{1,n}-\omega_{2,n})}\begin{pmatrix}\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})+\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\\[2.0pt] \big\{\omega_{1,n}\left[\varPhi_{1}(k_{n})+\varPhi_{1}(-k_{n})\right]+\omega_{2,n}\left[\varPhi_{2}(k_{n})+\varPhi_{2}(-k_{n})\right]\big\}/\mathrm{i}k_{n}\end{pmatrix}$,
where $\omega_{m,n}\equiv\omega_{m}(k_{n})$. The functions $\varPhi_{1,2}$ are given by Eq. (3.17), hence
[…]
@@
[…]
@@
$\boldsymbol{u}(x,t)\equiv\begin{pmatrix}e(x,t)\\[2.0pt] q(x,t)\end{pmatrix}=\begin{pmatrix}\varphi_{0}\\[3.0pt] 0\end{pmatrix}+\sum_{n=1}^{\infty}\frac{1}{\omega_{1,n}-\omega_{2,n}}\times\\[4.0pt] \times\begin{pmatrix}\big[(\omega_{1,n}\mathop{\mathrm{e}}\nolimits^{-\omega_{2,n}t}-\omega_{2,n}\mathop{\mathrm{e}}\nolimits^{-\omega_{1,n}t})\varphi_{n}+k_{n}(\mathop{\mathrm{e}}\nolimits^{-\omega_{1,n}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2,n}t})\psi_{n}\big]\cos(k_{n}x)\\[4.0pt] \big[-k_{n}\beta(\mathop{\mathrm{e}}\nolimits^{-\omega_{1,n}t}-\mathop{\mathrm{e}}\nolimits^{-\omega_{2,n}t})\varphi_{n}+(\omega_{1,n}\mathop{\mathrm{e}}\nolimits^{-\omega_{1,n}t}-\omega_{2,n}\mathop{\mathrm{e}}\nolimits^{-\omega_{2,n}t})\psi_{n}\big]\sin(k_{n}x)\end{pmatrix}$,
where $k_{n}=\pi n/l$, $\omega_{m,n}\equiv\omega_{m}(k_{n})$ (see Eq. (3.1)). This solution can also be easily obtained by the Fourier method.